Calculate the rate of heat conduction through house walls that are 10.5 cm thick and that have an average thermal conductivity twice that of glass wool. Assume there are no windows or doors. The surface area of the walls is 180 m2 and their inside surface is at 10.25 ºC, while their outside surface is at 4.23ºC


1
Expert's answer
2021-03-29T09:06:59-0400
"P=\frac{kA(T_h-T_c)}{d}\\
P=\frac{(0.084)(180)(10.25-4.23)}{0.105}=867\ W"